OpenStudy (anonymous):

let f(x)={x x<0 x^2 x>0 then f(2)-f(-2) equals A.2 B.-4 C.6 D.4 E.0

OpenStudy (shamim):

2>0 right?

OpenStudy (anonymous):

eyea

OpenStudy (anonymous):

yea

OpenStudy (shamim):

f(x)=x^2 f(2)=2^2 right?

OpenStudy (anonymous):

yes

OpenStudy (shamim):

wt do u think abt f(-2)